A boy and his sister stand on a balcony 14.0 m above the ground. The boy throws a ball velocity downward with an initial speed of 4.00 m/s. At the same instant, his sister throws a ball velocity upward with the same initial speed. When the first ball hits the ground, what is the vertical separation between the two balls.
1
Expert's answer
2020-08-21T10:44:51-0400

The law of motion of the "boy's ball"

"y_1=14.0-4.00t-\frac{9.81t^2}{2}"


The law of motion of the "girl's ball"

"y_2=14.0+4.00t-\frac{9.81t^2}{2}"

The time of motion of the first ball is given by equation

"14.0-4.00t-\frac{9.81t^2}{2}=0""t=1.33\:\rm s"

The position of the second ball at this instant

"y_2=14.0+4.00\times 1.33-\frac{9.81\times (1.33)^2}{2}=10.6\: \rm m"

Answer: "10.6\:\rm m"